### v1

@@ Abstract
Contrastive representation learning is crucial in time series analysis as it alleviates the issue of data noise and incompleteness as well as sparsity of supervision signal. However, existing constrastive learning frameworks usually focus on intral-temporal features, which fails to fully exploit the intricate nature of time series data. To address this issue, we propose DE-TSMCL, an innovative distillation enhanced framework for long sequence time series forecasting. Specifically, we design a learnable data augmentation mechanism which adaptively learns whether to mask a timestamp to obtain optimized sub-sequences. Then, we propose a contrastive learning task with momentum update to explore inter-sample and intra-temporal correlations of time series to learn the underlying structure feature on the unlabeled time series. Meanwhile, we design a supervised task to learn more robust representations and facilitate the contrastive learning process. Finally, we jointly optimize the above two tasks. By developing model loss from multiple tasks, we can learn effective representations for downstream forecasting task. Extensive experiments, in comparison with state-of-the-arts, well demonstrate the effectiveness of DE-TSMCL, where the maximum improvement can reach to 27.3%.
@@ 1 Introduction
Time series forecasting plays a critical role in various domains, ranging from finance and economics to weather forecasting and resource management[17, 23, 25]. The ability to accurately predict future values based on historical data is invaluable for decision-making and planning purposes. For the current time series forecasting study[9, 27, 37], the critical problem is how to leverage the inherent structure information of time series to learn discriminate representations to achieve better forecasting accuracy. In this article, the focus of our work is to efficiently integrate representation learning and forecasting into a unified framework and make them mutually supportive. With the emergence of contrastive learning methods, self-supervised representation learning and time series prediction have witnessed increased attention and progress. Several notable methods such as TS2Vec[37], TS-TCC [7], TF-C [40], and CPD [6] have been developed in this domain. The objective of contrastive learning is to train one or multiple encoders that can generate representations where similar instances are closer to each other, while dissimilar instances are farther apart. In particular, Yue et al. [37] introduced a comprehensive framework called TS2Vec for learning time series representations. This framework utilizes a hierarchical approach to distinguish between positive and negative samples. Eldele et al. [7] proposed a time-series representation learning framework called TS-TCC, which employs temporal and contextual contrastive learning to extract discriminative representations. Another notable contrastive learning approach, CoST, was introduced by Woo et al. [31] specifically for predicting long sequence time series. It transforms the data into the frequency domain to reveal the seasonal representations of the sequence, enabling a comprehensive understanding of its characteristics. These contrastive learning methods have demonstrated their effectiveness in learning informative representations for time series data.
Despite the promising results achieved in time series forecasting modeling, there are three crucial aspects that have often been neglected. These aspects have a significant impact on the accuracy and generalization ability of the models in real-world scenarios:
  • Data Noise and Distribution Shift: Real-world time series data often suffer from noise and incompleteness due to various environmental factors. This introduces additional challenges in accurate forecasting as the noise can corrupt the underlying patterns and relationships within the data. Moreover, the distribution of real-world data may differ from the training data, leading to a distribution shift. This inconsistency can result in poor model performance and inaccurate predictions.
@@
To address the above challenges, we develop DE-TSMCL, a distillation enhanced framework for time series forecasting based on momentum contrastive learning, which contains three key components: learnable data augmentation, distillation enhanced representation and momentum constrastive learning. First, we propose learnable data augmentation to learn whether to mask a sample to transform the original two overlapping subseries into enhanced views, which will be fed into the following teacher and student network and jointly optimized with the downstream forecasting task in an end-to-end fashion. Second, we introduce knowledge distillation technique into representation. We propose to train two models, teacher and student. Two models are trained simultaneously. The teacher make full use of of all available knowledge to capture the temporal features of time series to achieve better representation performance. Third, in order to calculate the similarity of samples from different timestamp, supervised task is considered. At the same time, considering the commonalities between multiple samples and effectively alleviating the problems of data sparsity, we design a self-supervised strategy powered with momentum contrastive training to further boost the performance. Finally, we jointly optimize these two tasks.
@@ 2 Related Work
In this section, we briefly review some related work on time series forecasting, contrastive learning and knowledge distilling.
@@ 2.1 Time Series Forecasting
In recent years, deep learning models have gained popularity in time series forecasting due to their ability to automatically learn complex patterns and dependencies. Convolution Neural networks(CNNS)[28, 1], Recurrent Neural Networks (RNNs)[30, 22, 26], particularly variants like Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU), have shown success in capturing long-term dependencies in time series data. Temporal Convolution Network (TCN) introduces dilated convolutions for time series forecasting and demonstrates their superior efficiency and predictive performance compared to RNNs[1]. Moreover, LSTnet [15] integrates CNNs and RNNs to simultaneously capture short-term local dependencies and long-term trends.
Additionally, attention mechanisms have been integrated into RNN-based models to focus on informative time steps or features, improving the model’s forecasting accuracy[16, 42, 32, 43]. Transformer-based architectures, originally designed for natural language processing tasks, have also been adapted for time series forecasting by employing self-attention mechanisms. LogTrans[16] employs convolutional self-attention layers with a LogSparse design to capture local information and reduce space complexity. Informer[42] introduces a ProbSparse self-attention mechanism that uses distilling techniques to efficiently extract the most crucial keys. Autoformer [32] incorporates the concepts of decomposition and auto-correlation from traditional time series analysis methods. FEDformer [43] employs a Fourier-transform structure to achieve linear complexity.
@@ 2.2 Contrastive Learning
In recent years, there has been growing interest in applying contrastive learning to time series forecasting. Oord et. al propose contrastive predictive coding (CPC)[21] to predict the subsequent latent variable, as opposed to negative samples drawn from the proposed distribution. Temporal and contextual contrasting (TS-TCC)[7], a variant of CPC[21], aims to optimize the concordance between robust and subtle enhancements of the same instance within an autoregressive framework. TS2Vec[37] treats enhanced views from the same time step as positive, while views from different time steps are considered negative. It also introduces instance-wise contrast between samples within the same batch. TNC[27] uses a discriminator network to predict subsequent data points. BTSF[35] creates positive pairs by applying a dropout layer to the same sample twice to minimize triplet loss associated with temporal and spectral features. Time-frequency consistency (TF-C)[40] is proposed to optimize the alignment between the temporal and frequency representations of the same instance. In addition, CoST[31] uses temporal consistencies in the time domain to learn the discriminative trend of a sequence, while transforming the data into the frequency domain to reveal the seasonal representation of the sequence.
2.3 Knowledge Distillition
@@
The knowledge distillation approach has been successfully applied in several domains, including time series forecasting [14]. Knowledge distillation is a technique in which a smaller model, called the student model, is trained to mimic the behavior and predictions of a larger, more complex model, called the teacher model. Existing knowledge distillation methods are classified into offline distillation [14] , online distillation [3, 2], and self-distillation [20]. Based on the type of knowledge, current online KD approaches can be mainly categorized into response-based[41], feature-based[5], and relation-based methods[34]. Fan et al.[8] randomly select two subsequences of length $L$ from the same time series and assign pseudo-labels based on their temporal separation. Then, the proposed model is pre-trained to predict the pseudo-labels of the subsequence pairs. Zhang et al.[39] have integrated expert features to generate pseudo-labels for self-supervised contrastive representation learning of time series. Conventional models train and test on identical data, missing information after prediction and before the result. LuPIET[18] includes this missing data in training to improve predictions by distilling knowledge. However, the prediction of pseudo-labels is prone to include incorrect labels. Therefore, a key focus in the study of training is to mitigate the negative effects of these incorrect labels.
@@ 3.1 Problem Statement
Let $\mathcal{X}=\{\mathbf{x_{1}},\mathbf{x_{2}},...,\mathbf{x_{N}}\}\in\mathbb{R}^{N\times T\times C}$ represent a set of time series, where $N$ is the number of instances. Each vector $\mathbf{x_{i}}$ has dimension $T\times C$ , where $T$ is the length of time series and $C$ represents the number of channels or features. Given the the lookback window $T$, the time series forecasting problem aims to predict the time series at the next $P$ steps, $\hat{\mathbf{x}}=f(\mathbf{x})$, where $\mathbf{x}\in\mathbb{T}^{T\times C}$ is the historical observations, and $\hat{\mathbf{x}}\in\mathbb{R}^{P\times C}$ is the future $P$ steps predictions,
Figure 1: The overall architecture of DE-TSMCL. It consists of four major components: data augmentation and representation learning, supervised task, and self-supervised task.
@@ 3.2 Momentum Contrastive Learning
Consider a sample query $q$ generated by a query encoder and a sequence of samples, denoted $k=\{k_{1},k_{2},...,k_{N}\}$, generated by a momentum encoder. Suppose the unique pairs of $k$ and $q$ in our sample are extremely similar, having been generated by the same encoder with similar parameters. In this case, we should minimize the distance between the most similar pair $q$ and $k^{+}$ and increase the distance between other pairs that are not so similar $q$ and other $k^{-}$.
@@
This method quantifies similarity using the dot product. Previous work has called it InfoNCE, which is a variant of the NCE loss function first proposed in CPC[21], and since then further developed for applications such as MOCO [11] and GCA[36].
@@ 4.1 Overall Framework
Our approach follows the teacher-student training scheme. Both the student and teacher networks share the same underlying architecture but possess distinct parameters In each training iteration, we first randomly sample the original time series $\mathbf{x}$ to form two overlapping sub-series, namely $\mathbf{x_{s}}$ and $\mathbf{x_{t}}$, and perform data augmentation on each sub-series to obtain the augmented view, which are then fed into the teacher and student network, respectively, to generate representation $\mathbf{h_{s}}$ and $\mathbf{h_{t}}$. $\mathbf{h_{t}}$ is utilized as supervision for the teacher network. Next, we set up two tasks: a self-supervise task and an adaptive supervised task. In self-supervised task, we conduct momentum contrastive learning on $\mathbf{h_{s}}$ and $\mathbf{h_{t}}$ to capture the complex temporal dependencies among time series and alleviate data noise and incompleteness. In supervised task, we employ a conventional cross-entropy loss to more effectively align the semantic information at the same timestamp, thereby facilitating better updated model parameters through back-propagation. Finally, we jointly optimize the two tasks.
@@
$\mathcal{X}\in\mathbb{R}^{N\times T\times C}$
Time series set. $N$ is the number of instances, $T$ is
the length of lookback window, $C$ is the number of
features.
$\mathcal{Z}\in R^{N\times T\times C}$  The set of latent vector after Projection(MLP).
$\mathcal{H}\in R^{N\times L\times K}$
The learned instance representations set. $L$ is the
length of overlapping subseries, $K$ is the dimension
of representations.
@@
$z_{i}$
The latent vector after Projection(MLP) of instance
$i$.
@@
The size of Batchsize and overlapping subsequence
Length.
@@
The parameters sets of teacher and student
encoders.
@@
The weight for balance loss, and the momentum
coefficient.
@@
Element-wise addition / minus / multiplication /
division.
@@ 4.2 Data Processing and Augmentation
In order to conserve the macro patterns of time series and facilitate the encoder effectively extract the high-level information, we design the data augmentation block. The significance of data augmentation in contrastive learning has been acknowledged in previous studies[4, 10], data augmentation plays an important role in contrastive learning. While existing models commonly employ random masking techniques to refine instance representations $\mathbf{x}$, such an approach often introduces biased and noisy information. Moreover, relying solely on masking mechanisms in contrastive learning for time series forecasting is insufficient for generating powerful representations capable of mitigating the aforementioned biases and noises. To address these limitations, we propose the utilization of parameterized networks for generating optimized representations. Furthermore, the quadratic increase in memory and computational costs associated with the augmentation of multiple views poses practical challenges. To tackle this issue, we introduce a dual-cropping strategy wherein two overlapping sub-series are randomly sampled. In addition, we incorporate learnable augmentation strategies to enhance the robustness of the learned representations.
@@ 4.2.1 Dual-cropping
We denote the time series $\mathcal{X}=\{\mathbf{x_{i}}\}_{i=1}^{N}$ and the corresponding representations $H=\{\mathbf{h_{i}}\}_{i=1}^{N}$. Firstly, for each instance $\mathbf{x_{i}}$, we first sample two overlapping sub-series, denoted as $\mathbf{x_{i}^{1}}=\{x_{i,t}\}_{t=a_{1}}^{b_{1}}$ and $\mathbf{x_{i}^{2}}=\{x_{i,t}\}_{t=a_{2}}^{b_{2}}$, where $\{a_{1},b_{1},a_{2},b_{2}\}\in\{1,2,\cdots,N\}$, satisfying the conditions $a_{1}<b_{1}$, $b_{1}<a_{2}$, and $a_{2}<b_{2}$. The selection of two sub-sequences from the same time series is crucial as they contain identical semantic information. Consequently, these two subsequences can be considered as positive sample pairs. This approach, pioneered by Word2vec [19], has been widely adopted in various subseries studies. Without loss of generality, for a given instance $\mathbf{x}_{i}$, we randomly choose a value $0<L\leq T$ and subsequently extract two overlapping subsequences $\mathbf{x_{i}^{1}}$ and $\mathbf{x_{i}^{2}}$ of equal window size $L$.

### v2

@@ Abstract
Contrastive representation learning is crucial in time series analysis as it alleviates the issue of data noise and incompleteness as well as sparsity of supervision signal. However, existing constrastive learning frameworks usually focus on intral-temporal features, which fails to fully exploit the intricate nature of time series data. To address this issue, we propose DE-TSMCL, an innovative distillation enhanced framework for long sequence time series forecasting. Specifically, we design a learnable data augmentation mechanism which adaptively learns whether to mask a timestamp to obtain optimized sub-sequences. Then, we propose a contrastive learning task with momentum update to explore inter-sample and intra-temporal correlations of time series to learn the underlying structure feature on the unlabeled time series. Meanwhile, we design a supervised task to learn more robust representations and facilitate the contrastive learning process. Finally, we jointly optimize the above two tasks. By developing model loss from multiple tasks, we can learn effective representations for downstream forecasting task. Extensive experiments, in comparison with state-of-the-arts, well demonstrate the effectiveness of DE-TSMCL, where the maximum improvement can reach to 27.3%. Source code for the algorithm is available at https://github.com/gaohaozhi/DE-TSMCL.
@@ 1 Introduction
Time series forecasting plays a critical role in various domains, including finance, economics, weather forecasting, and resource management. Accurately predicting future values based on historical data is essential for decision-making and planning purposes in these fields[19, 24, 26]. For the current time series forecasting study, the critical problem is how to leverage the inherent structure information of time series to learn discriminate representations to achieve better forecasting accuracy[29, 42]. The ability of a model to learn these representations is critical for enhancing its performance[37, 10]. The primary focus of our work in this article is to develop an efficient framework that integrates representation learning and forecasting in a unified manner.
With the emergence of contrastive learning methods, self-supervised representation learning and time series prediction have witnessed increased attention and progress. Several notable methods such as TS2Vec[42], TS-TCC [8], TF-C [45], and CPD [7] have been developed in this domain. The objective of contrastive learning is to train one or multiple encoders that can generate representations where similar instances are closer to each other, while dissimilar instances are farther apart. In particular, Yue et al. [42] introduce a comprehensive framework called TS2Vec for learning time series representations. This framework utilizes a hierarchical approach to distinguish between positive and negative samples. Eldele et al. [8] propose a time-series representation learning framework called TS-TCC, which employs temporal and contextual contrastive learning to extract discriminative representations. Another notable contrastive learning approach, CoST, is introduced by Woo et al. [33] specifically for predicting long sequence time series. It transforms the data into the frequency domain to reveal the seasonal representations of the sequence, enabling a comprehensive understanding of its characteristics. These contrastive learning methods have demonstrated their effectiveness in learning informative representations for time series data.
Despite the promising results in time series forecasting, there are three crucial aspects that have often been neglected. These aspects have a significant impact on the accuracy and generalization ability of the models in real-world scenarios:
  • Data Noise and Distribution Shift: Real-world time series data often suffers from noise and incompleteness due to various environmental factors. This introduces additional challenges in accurate forecasting as the noise can corrupt the underlying patterns and relationships within the data. Moreover, the distribution of real-world data may differ from the training data, leading to a distribution shift. This inconsistency can result in poor model performance and inaccurate predictions.
@@
To address the above challenges, we develop DE-TSMCL, a distillation enhanced framework for time series forecasting based on momentum contrastive learning, which contains three key components: learnable data augmentation, distillation enhanced representation and momentum constrastive learning. First, we propose learnable data augmentation to learn whether to mask a sample to transform the original two overlapping subseries into enhanced views, which will be fed into the following teacher and student network and jointly optimized with the downstream forecasting task in an end-to-end fashion. Second, we introduce knowledge distillation technique into representation learning process. We simultaneously train two models, the teacher and the student. The teacher makes full use of of all available knowledge to capture the temporal features of time series to achieve better representation performance. Third, in order to calculate the similarity of samples from different timestamp, supervised task is considered. At the same time, considering the commonalities between multiple samples and effectively alleviating the problems of data sparsity, we design a self-supervised strategy powered with momentum contrastive training to further boost the performance. Finally, we jointly optimize these two tasks.
@@ 2 Related Work
In this section, we briefly review some related work on time series forecasting, contrastive learning and knowledge distillation.
@@ 2.1 Time Series Forecasting
In recent years, deep learning models have been widely used in time series forecasting due to their ability to automatically learn complex patterns and dependencies. Convolution Neural networks(CNNs)[30, 1], Recurrent Neural Networks (RNNs)[32, 23, 27], particularly variants like Long Short-Term Memory (LSTM) and Gated Recurrent Units (GRU), have shown success in capturing long-term dependencies in time series data. Temporal Convolution Network (TCN[1]) introduces dilated convolutions for time series forecasting and demonstrates their superior efficiency and prediction performance compared to RNNs[1]. Moreover, LSTNet [17] integrates CNNs and RNNs to simultaneously capture short- and long-term dependencies.
In addition, attention mechanisms have been integrated into RNN-based models to improve the model’s forecasting accuracy[18, 47, 34, 48]. Transformer-based architectures, originally designed for natural language processing tasks, have also been adapted for time series forecasting by employing self-attention mechanisms. LogTrans[18] employs convolutional self-attention layers with a LogSparse design to capture local information features. Informer[47] introduces a ProbSparse self-attention mechanism that uses distillation techniques to efficiently extract the most crucial keys. Autoformer [34] incorporates the concepts of decomposition and auto-correlation from traditional time series analysis methods. FEDformer [48] employs a Fourier-transform structure and achieve linear complexity.
@@ 2.2 Contrastive Learning
In recent years, there has been growing interest in applying contrastive learning for time series forecasting. Oord et. al propose contrastive predictive coding (CPC)[22] to predict the subsequent latent variable, as opposed to negative samples drawn from the proposed distribution. Temporal and contextual contrasting (TS-TCC)[8], a variant of CPC[22], aims to optimize the concordance between robust and subtle enhancements of the same instance within an autoregressive framework. TS2Vec[42] treats enhanced views from the same time step as positive, while views from different time steps as negative. It also introduces instance-wise contrast between samples within the same batch. TNC[29] uses a discriminator network to predict subsequent data points. BTSF[40] creates positive pairs by applying a dropout layer to the same sample twice to minimize triplet loss associated with temporal and spectral features. TF-C[45] is proposed to optimize the alignment between the temporal and frequency representations of the same instance. In addition, CoST[33] uses temporal consistencies in the time domain to learn the discriminative trend of a sequence, while transforming the data into the frequency domain to reveal the seasonal representation of the sequence.
2.3 Knowledge Distillation
@@
The knowledge distillation approach has been successfully applied in several domains, including time series forecasting [15]. Knowledge distillation is a technique in which a smaller model, called the student model, is trained to mimic the behavior and predictions of a larger and more complex model, called the teacher model. Existing knowledge distillation methods are classified into offline distillation [15] , online distillation [4, 3], and self-distillation [21]. Based on the type of knowledge, current online knowledge distillation approaches can be mainly categorized into response-based[46], feature-based[6], and relation-based methods[39]. Fan et al.[9] randomly select two subsequences from the same time series and assign pseudo-labels based on their temporal separation. Then, the proposed model is pre-trained to predict the pseudo-labels of the subsequence pairs. Zhang et al.[44] integrate expert features to generate pseudo-labels for self-supervised contrastive representation learning of time series. LuPIET[20] considers missing data in training to improve predictions with distillation knowledge. However, the prediction of pseudo-labels is prone to include incorrect labels. Therefore, a key focus in the study of training is to mitigate the negative effects of these incorrect labels.
Recently, some works have exploited the self-distillation framework to achieve good results on time series classification, and other classification and recognition tasks[38, 36]. In particular, Xiao et.al [38] combines data augmentation, deep contrastive learning, and self-distillation. It considers the contrast similarity of both the high- and low-level semantic information. CapMatch [36] focuses on extracting rich representations from input data using a hybrid approach that combines supervised and unsupervised learning techniques. In addition, it construct similarity learning on lower and higher-level semantic information extracted from augmented data to recognize correlations. Different from existing studies, our work focuses on the integration of representation learning and forecasting into a unified framework. Furthermore, we introduce knowledge distillation between teacher and student models to improve representation performance. We also address the similarity of samples from different timestamps using supervised tasks and incorporate a self-supervised strategy to alleviate data sparsity issues.
@@ 3.1 Problem Statement
Let $X=\{\mathbf{x_{1}},\mathbf{x_{2}},...,\mathbf{x_{N}}\}\in{R}^{N\times T\times C}$ represent a set of time series, where $N$ is the number of instances. Each vector $\mathbf{x_{i}}$ has dimension $T\times C$ , where $T$ is the length of time series and $C$ represents the number of channels or features. Given the the look-back window $T$, the time series forecasting problem aims to predict the time series at the next $P$ steps, $\hat{\mathbf{x}}=f(\mathbf{x})$, where $\mathbf{x}\in R^{T\times C}$ is the historical observations, and $\hat{\mathbf{x}}\in R^{P\times C}$ is the future $P$ steps predictions,
Figure 1: The overall architecture of DE-TSMCL. It consists of four major components: data augmentation, representation learning, supervised task, and self-supervised task.
@@ 3.2 Momentum Contrastive Learning
Consider a sample query $q$ generated by a query encoder and a sequence of samples, denoted $k=\{k_{1},k_{2},...,k_{N}\}$, generated by a momentum encoder. Suppose the unique pairs of $k$ and $q$ in our sample are extremely similar, having been generated by the same encoder with similar parameters. In this case, the objective is to minimize the distance between the most similar pair, represented as $q$ and $k^{+}$, while increasing the distance between other pairs where the similarity between $q$ and other $k^{-}$ is not as significant.
@@
Contrastive learning measures similarity using the dot product. It has been previously referred to as InfoNCE, which is a variation of the NCE loss function initially introduced in CPC[22]. Since then, it has been further developed and applied in various contexts such as MOCO[12] and GCA[41].
@@ 4.1 Overall Framework
Our approach follows the teacher-student training scheme. Both the student and teacher networks share the same underlying architecture but possess distinct parameters. In each training iteration, we first randomly sample the original time series $\mathbf{x}$ to form two overlapping sub-series, namely $\mathbf{x_{s}}$ and $\mathbf{x_{t}}$, and perform data augmentation on each sub-series to obtain the augmented view, which are then fed into the teacher and student network, respectively, to generate representation $\mathbf{h_{s}}$ and $\mathbf{h_{t}}$. $\mathbf{h_{t}}$ is utilized as supervision for the teacher network. Next, we set up two tasks: a self-supervise task and an adaptive supervised task. In self-supervised task, we conduct momentum contrastive learning on $\mathbf{h_{s}}$ and $\mathbf{h_{t}}$ to capture the complex temporal dependencies among time series and alleviate data noise and incompleteness. In supervised task, we employ a conventional cross-entropy loss to more effectively align the semantic information at the same timestamp, thereby facilitating better updated model parameters through back-propagation. Finally, we jointly optimize the two tasks.
@@
${X}\in{R}^{N\times T\times C}$
Time series set. $N$ is the number of instances, $T$ is the length of look-back window,
$C$ is the number of features.
${H}\in R^{N\times L\times K}$
The learned instance representations set. $L$ is the length of overlapping subseries,
$K$ is the dimension of representations.
@@
$\mathbf{z}_{\mathbf{i}}\in R^{T\times C}$
The latent vector after MLP of instance $i$.
@@
The size of batchsize and overlapping sub-series length.
@@
The parameters sets of teacher and student encoders.
@@
The weight for balance loss, and the momentum coefficient.
@@
Element-wise addition / minus / multiplication / division.
@@ 4.2 Data Processing and Augmentation
In order to conserve the macro patterns of time series and facilitate the encoder effectively extract the high-level information, we design the data augmentation block. The significance of data augmentation in contrastive learning has been acknowledged in previous studies[5, 11]. While existing models commonly employ random masking techniques to refine instance representations, such an approach often introduces biased and noisy information. Moreover, relying solely on masking mechanisms in contrastive learning for time series forecasting is insufficient for generating powerful representations capable of mitigating the aforementioned biases and noises. To address these limitations, we propose the utilization of parameterized networks for generating optimized representations. Furthermore, the quadratic increase in memory and computational costs associated with the augmentation of multiple views poses practical challenges. To tackle this issue, we introduce a dual-cropping strategy wherein two overlapping sub-series are randomly sampled. In addition, we incorporate learnable augmentation strategies to enhance the robustness of the learned representations.
@@ 4.2.1 Dual-cropping
We denote the time series $X=\{\mathbf{x_{i}}\}_{i=1}^{N}$ and the corresponding representations $H=\{\mathbf{h_{i}}\}_{i=1}^{N}$. Firstly, for each instance $\mathbf{x_{i}}$, we first sample two overlapping sub-series, denoted as $\mathbf{x_{i}^{1}}=\{x_{i,t}\}_{t=a_{1}}^{b_{1}}$ and $\mathbf{x_{i}^{2}}=\{x_{i,t}\}_{t=a_{2}}^{b_{2}}$, where $\{a_{1},b_{1},a_{2},b_{2}\}\in\{1,2,\cdots,N\}$, satisfying the conditions $a_{1}<b_{1}$, $b_{1}>a_{2}$, and $a_{2}<b_{2}$. The selection of two sub-sequences from the same time series is crucial as they contain identical semantic information. Without loss of generality, for a given instance $\mathbf{x_{i}}$, we randomly choose a value $0<L\leq T$ and subsequently extract two overlapping subsequences $\mathbf{x_{i}^{1}}$ and $\mathbf{x_{i}^{2}}$ of equal window size $T$, and the length of the overlap between two sub-series is $L$.